# Emergence of a Landau level structure in dark optical lattices

### v2

@@ Abstract
An optical flux lattice is a set of light beams that couple different internal states of an atom, thereby producing topological energy bands. Here we present a configuration in which the atoms exhibit a dark state, i.e. an internal state that is not coupled to the light. At large light intensity, the low-energy dynamics is restricted to the dark state, leading to an effective continuum model with a Landau-level-like structure. This structure is dramatically different from that of usual topological optical lattices, which lead to discrete models in the tight-binding limit. The proposed system is essentially immune to heating due to photon scattering, making it a highly promising way to emulate the integer or fractional quantum Hall effect.
@@
The generation of topological bands for atoms or photons moving in a periodic potential has been recently the subject of many theoretical and experimental studies (for reviews, see e.g. [1, 2]). It opens the way to a simulation of the quantum Hall effect, from the integer to the fractional regimes when the role of interactions increases. Among the various schemes that have been considered for atoms moving in a laser field, optical flux lattices (OFL) have the advantage of not requiring any time modulated parameters (which may lead to undesired heating), while providing a large flux density, typically one or two flux quanta per unit cell [3]. However, OFL tend to require relatively complex laser beam structures, when they are aimed at atomic species, such as alkali-metal atoms, commonly used in quantum gas experiments [4, 5]. In addition, OFL schemes proposed so far may induce significant photon scattering, which limits the available time during which the atomic gas remains cold enough to observe topology-induced effects.
@@
In this paper, we propose a modified version of the optical lattice concept that operates with a so-called ‘dark state’, thus reducing considerably the heating problem due to photon scattering. Dark state lattices have recently been identified as a method for achieving sub-wavelength structures, either in 1D [6, 7, 8] or 2D [9]. Here we propose a scheme providing an energy spectrum remarkably close to the Landau level structure for a charged particle in a uniform magnetic field. For realistic parameters, we obtain a series of many equidistant, non-overlapping, topological bands, each with a Chern index equal to 1. Being restricted to the dark internal state, the atoms remain weakly coupled to light, giving rise to delocalized orbitals. This behavior contrasts with other realizations of topological bands in optical lattices, which can be described by discrete lattice models such as the Haldane and Harper-Hofstadter models [10, 11, 12]. As an illustration of the topological character of the bands, we check that an ideal Fermi gas populating the ground band should exhibit an incompressible bulk surrounded by chiral edge states, and interacting bosons should organize themselves into an ordered vortex lattice.
@@
We consider atoms with an electronic ground level $g$ with angular momentum $J_{g}=1$, moving in the $xy$ plane. The laser beams forming the OFL are supposed to be monochromatic and they couple the ground level $g$ to an excited level $e$, also with angular momentum $J_{e}=1$. One can then identify a space-dependent dark state, linear combination of $|g,m_{z}=\pm 1\rangle$, where $m_{z}$ labels the projection of the internal angular momentum along $z$ [13].
Since the polarization vector $\bm{\epsilon}(\bm{r})$ lies in the $xy$ plane, it can decomposed at any point $\bm{r}$ on the circular polarization basis ϵ±subscriptbold-italic-$\bm{\epsilon}_{\pm}$, $\bm{\epsilon}(\bm{r})\propto\alpha_{+}(\bm{r})\bm{\epsilon}_{+}\,+\,\alpha_{-}(\bm{r})\bm{\epsilon}_{-}$, where $\alpha_{\pm}(\bm{r})$ are complex, spatially periodic functions. The atom-laser coupling reads in the $|g,\pm 1\rangle$ basis:
$\hat{V}(\bm{r})=V_{0}\begin{pmatrix}|\alpha_{-}|^{2}&-\alpha_{+}\alpha_{-}^{*}\\ -\alpha_{+}^{*}\alpha_{-}&|\alpha_{+}|^{2}\end{pmatrix},$ (1)
where $V_{0}$ is a real, positive energy. Here we restricted the internal atomic dynamics to the $\Lambda$ system formed by $\{|g,\pm 1\rangle,|e,0\rangle\}$, and we adiabatically eliminated the excited state $|e,0\rangle$ assuming that the positive detuning from resonance is large compared to the Rabi frequency characterizing the atom-laser coupling (see appendix 1 for details) [14].
The dark state $|{\cal D}(\bm{r})\rangle\propto\alpha_{+}(\bm{r})|g,+1\rangle+\alpha_{-}(\bm{r})|g,-1\rangle$ is the eigenstate of $\hat{V}$ with energy 0, which provides a one-to-one correspondance between the Poincaré sphere for the polarization, and the Bloch sphere for the two-level system $|g,\pm 1\rangle$. The other eigenstate of $\hat{V}$ is the bright state $|{\cal B}(\bm{r})\rangle\propto\alpha_{-}^{*}(\bm{r})|g,+1\rangle-\alpha_{+}^{*}(\bm{r})|g,-1\rangle$. Its energy $E_{\text{bright}}({\bm{r}})=V_{0}(|\alpha_{+}({\bm{r}})|^{2}+|\alpha_{-}({\bm{r}})|^{2})$ is positive everywhere so that the dark state is always the local ground state of the atom.
In this work, we will choose the following expressions for the coefficients $\alpha_{\pm}(\bm{r})$  [^1] The electric field corresponding to eqs. (2-3) is not transverse and therefore cannot be implemented as such. When using a spatial light modulator in the paraxial regime, the generation of the amplitudes $\alpha_{\pm}(\bm{r})$ involves an additional polarisation component along the $z$ axis. However, this third component has a comparatively small amplitude and therefore does not change our conclusions. Alternatively, one can lift the degeneracy between the $\sigma_{\pm}$ transitions with an external static magnetic field and use the two-frequency configuration described at the end of Appendix 1 to achieve the desired amplitudes.:
$\alpha_{+}(\bm{r})$ $=$ $1+\frac{1}{2}\left[\cos(kx)+\cos(ky)\right],$ (2)
$\alpha_{-}(\bm{r})$ $=$ $\sin\left[\frac{k}{2}(x-y)\right]-{\rm i}\sin\left[\frac{k}{2}(x+y)\right].$ (3)
The light-shift coupling is invariant upon discrete symmetries that combine discrete translations by $d\,{\bm{e}}_{x}$ or $d\,{\bm{e}}_{y}$ and the spin rotation $\hat{\sigma}_{z}$, where $d=2\pi/k$, which allows us to define a square $d\times d$ unit cell.
The main results of this paper are based on the numerical calculation of the energy bands and of the associated Bloch states for a two-level atom moving in the potential $V(\bm{r})$. This will be done without any assumption regarding the adiabatic following of the dark state. However, it is interesting to start with the case when this adiabatic following is assumed. A first remark is that the dark state wraps the Bloch sphere, which is a necessary condition for the operation of an OFL. For example, the north pole $|g,+1\rangle$ is reached when ${\bm{r}}\to\bm{0}$, and the south pole $|g,-1\rangle$ when ${\bm{r}}\to(d/2,d/2)$. A consequence of this wrapping is that the flux across the unit cell of the synthetic magnetic field associated with the adiabatic following of the dark state $|{\cal D}(\bm{r})\rangle$ is nonzero. It is equal to $2\pi\hbar$, corresponding to one flux quantum across the unit cell.
This non-zero flux is one of the requirements for an OFL. Another requirement formulated in [3] is that the eigenstates of the coupling $V(\bm{r})$ should be non-degenerate at any point $\bm{r}$. This second requirement is clearly not fulfilled for the choice (2-3) since $\hat{V}$ vanishes when ${\bm{r}}\to(d/2,d/2)$. In fact, this second requirement cannot be satisfied for a dark-state based OFL, as shown in appendix 2. However, as we see below, this does not prevent us to reach well separated topological bands.
@@
Figure 1: (a) Band spectrum of the flux lattice as a function of the light coupling amplitude $V_{0}$. In the regime $V_{0}\gg\hbar\omega_{\text{c}}$, we find at low energy a succession of narrow energy bands with dominant dark-state character and an almost uniform spacing $\simeq\hbar\omega_{\text{c}}$. The red lines give a rough estimate of the first two bands with dominant bright-state character. They correspond to the energy levels $(n+1)\hbar\Omega_{\text{bright}}$ (with $n$ integer) in the lattice potential experienced by the bright state in the large depth limit of independent harmonic wells. (b) Mean value $\langle\hat{V}\rangle$ of the light shift as a function of the mean total energy $\langle E\rangle$, averaged over the quasi-momentum ${\bm{q}}$, for the successive bands. The small value of the light shift for the low-energy bands shows that they predominantly populate the dark state.
@@
We show in Fig. 1a the band spectrum for a two-level atom for the Hamiltonian $\hat{p}^{2}/(2m)+\hat{V}(\bm{r})$ as a function of the light coupling amplitude $V_{0}$. Here, energies are expressed in units of the cyclotron frequency $\hbar\omega_{\text{c}}=2\pi\hbar^{2}/(md^{2})$ expected for a Landau level with the same magnetic flux density as our OFL. For large couplings $V_{0}\gg\hbar\omega_{\text{c}}$, we find a series of narrow energy bands at low energy, with an almost uniform spacing matching the cyclotron gap $\hbar\omega_{\text{c}}$. These bands essentially populate the dark state, with a very small residual light shift, shown in Fig. 1b for $V_{0}=100\,\hbar\omega_{\text{c}}$. For this light coupling, the ground band exhibits a width $\simeq 0.11\,\hbar\omega_{\text{c}}$ (about 7.2 times smaller than the gap to the first excited band), and the mean light shift is as low as $\langle\hat{V}\rangle\simeq 10^{-4}V_{0}$. The mixing between dark and bright states occurs at a higher energy, which can be estimated by computing the spectrum of an atom subjected to the lattice potential given by the position-dependent bright-state energy $E_{\text{bright}}({\bm{r}})$. In the large depth limit, one can neglect tunneling between the different wells of the bright-state lattice, and one expects a harmonic spectrum for that state of frequency $\Omega_{\text{bright}}=\sqrt{2\pi V_{0}\omega_{\text{c}}/\hbar}$ (dashed red lines in Fig. 1a).
To characterize the topological nature of the bands, we compute their Chern number, and find the value $C=1$ for the first 19 bands at $V_{0}=100\,\hbar\omega_{\text{c}}$, confirming the similarity with Landau levels. Furthermore, the Berry curvature of the ground band is relatively uniform, with an rms deviation equal to $\simeq 25\%$ of its mean value.
Recent theoretical developements put forward a class of ‘ideal’ Chern bands, which share the same algebraic structure as Landau levels despite a non-uniform Berry curvature [16, 17, 18, 19]. These bands are characterized by a quantum geometrical tensor $[\mathcal{Q}_{\bm{q}}]$ exhibiting a constant null vector ${\bm{w}}$, i.e. $[\mathcal{Q}_{\bm{q}}]{\bm{w}}=\bm{0}$ for all ${\bm{q}}$ [17]. For the ground band of our OFL, we find that the circular vector ${\bm{w}}=\bm{\epsilon}_{-}$ acts as an null vector to a very good approximation (on average over the band, we find $||[\mathcal{Q}_{\bm{q}}]\bm{\epsilon}_{-}||=0.06||[\mathcal{Q}_{\bm{q}}]\bm{\epsilon}_{+}||$). Consequently, the dynamics of the momentum component $q_{x}-{\rm i}q_{y}$ is effectively suppressed, giving rise to momentum-space holomorphicity, akin to Landau levels.
@@
We can extend the comparison between the OFL and Landau levels by inspecting individual Bloch states in real space.
We first recall the Landau level formalism in a way that is well suited for comparison with flux lattices, i.e. a spatially periodic problem. We focus here on the ground band, the lowest Landau level (LLL). We assume a magnetic field ${\bm{B}}=B\,{\bm{e}}_{z}$ and adopt the Landau gauge ${\bm{A}}=-By\,{\bm{e}}_{x}$, for which the system is invariant upon $x$-translations $t_{x}(d_{x})=\exp(-d_{x}\,\partial_{x})$ of arbitrary distance $d_{x}$. The resulting conservation of the canonical momentum $p_{x}$ yields the LLL basis states $\lvert\phi_{p_{x}}\rangle$, of wavefunction
@@
where $\ell=\sqrt{\hbar/qB}$ is the magnetic length. The basis of $\lvert\phi_{p_{x}}\rangle$ is not convenient for comparison with OFL Bloch bands, which are indexed by a 2D quasi-momentum ${\bm{q}}$. In order to symmetrize the $x$ and $y$ directions, we use another symmetry of the system, namely the $y$-magnetic translation symmetry $t_{y}(d_{y})=\exp[-d_{y}\,(\partial_{y}-{\rm i}x/\ell^{2})]$ [20]. The commutation of translation symmetries obeys the algebra $t_{x}(d_{x})t_{y}(d_{y})={\rm e}^{-{\rm i}d_{x}d_{y}/\ell^{2}}t_{y}(d_{x})t_{x}(d_{y})$, hence a common basis of $x$- and $y$-translations can be found by choosing $d_{x}=d_{y}=d=\sqrt{2\pi}\ell$, such that the $d\times$d unit cell is thread by a single unit of magnetic flux quantum $\Phi_{0}=h/q$. Magnetic Bloch states $\lvert\psi_{q_{x},q_{y}}\rangle$ invariant under both $t_{x}(d)$ and $t_{y}(d)$ can be formed by discrete sum of momentum states $\lvert\phi_{p_{x}}\rangle$, as
$\lvert\psi_{q_{x},q_{y}}\rangle=\sqrt{d}\sum_{n=-\infty}^{\infty}{\rm e}^{{\rm i}q_{y}dn}\lvert\phi_{q_{x}+nQ}\rangle.$ (5)
This expression corresponds to a discrete Fourier transform of all the momentum states $\lvert\phi_{q_{x}+nQ}\rangle$ whose density probabilities are shifted one to the other along $y$ by integer multiples of $d$. The states $\lvert\psi_{q_{x},q_{y}}\rangle$, defined in the Brillouin zone $|q_{x}|,|q_{y}|<\pi/d$, form a basis of the LLL. We note that the wavefunctions match the well-known LLL states defined on a torus of area $d\times d$ thread by magnetic fluxes $2\pi(\Phi_{x},\Phi_{y})/\Phi_{0}=(q_{x},q_{y})d$, given by [21]
$\psi_{q_{x},q_{y}}(x,y)=\frac{{\rm e}^{-y^{2}/(2\ell^{2})-{\rm i}q_{x}q_{y}d^{2}/(2\pi)}}{\pi^{1/4}\sqrt{\ell}}\Theta\begin{bmatrix}q_{x}d\\ q_{y}d\end{bmatrix}\left(\frac{x-{\rm i}y}{d}\middle|{\rm i}\right),$ (6)
@@
Figure 2: Mean density (a,c) and velocity (b,d) distributions on a square unit cell $d\times d$ thread by one unit of magnetic flux quantum. We assume a Bloch state ${\bm{q}}=\bm{0}$, which exhibits a quantized vortex at $(d/2,d/2)$. The left (resp. right) panels correspond to the lowest band of the OFL for $V_{0}=100\,\hbar\omega_{\text{c}}$ (resp. the LLL). In (b,d), the color encodes the norm of the velocity, and the red arrows the velocity streamlines.
@@
is shown in Fig. 2d. We note that the circulation of velocity around each unit cell cancels due to the compensation between the singular contribution of the vortex and the rotational flow induced by the magnetic field. Each magnetic Bloch state thus exhibits non-zero local currents within the magnetic unit cell, but no global flow on larger length scales. The occurrence of vortices in OFL orbitals provides another illustration of the ‘ideal’ Chern band character: like in the LLL, OFL orbitals can be fully characterized by the location of quantum vortices [18].
Figure 3: In situ density (a) and current density (b) profile of a Fermi gas at zero temperature in a quasi-2D geometry subjected to the optical flux lattice with $V_{0}=100\,\hbar\omega_{\text{c}}$ and to an isotropic harmonic trap (trapping frequency $\omega_{\text{trap}}=0.025\,\omega_{\text{c}}$). We used a chemical potential $\mu_{0}=1.11\,\hbar\omega_{\text{c}}$ located between the ground and first excited bands. A few stream lines are shown as red lines in (b). (c) Density profile along the line $y=0$, showing incompressibility in the bulk.

### v3

@@ Abstract
An optical flux lattice is a set of light beams that couple different internal states of an atom, thereby producing topological energy bands. Here we present a configuration in which the atoms exhibit a dark state, i.e. an internal state that is not coupled to the light. At large light intensity, the low-energy dynamics is restricted to the dark state, leading to an effective continuum model with a Landau-level-like structure. This structure is dramatically different from that of usual topological optical lattices, which lead to discrete models in the tight-binding limit. For well-chosen atomic species, the proposed system is essentially immune to heating due to photon scattering, making it a highly promising way to emulate the integer or fractional quantum Hall effect.
@@
The generation of topological bands for atoms or photons moving in a periodic potential has been recently the subject of many theoretical and experimental studies (for reviews, see e.g. [1, 2]). It opens the way to a simulation of the quantum Hall effect, from the integer to the fractional regimes when the role of interactions increases. Among the various schemes that have been considered for atoms, optical flux lattices (OFL) involve laser fields inducing a spatially periodic coupling between different internal states of the atom. Suitable laser configurations give rise to topological Bloch bands, with a periodic effective magnetic flux allowing cold atoms to experience conditions analogous to those found in quantum Hall states. OFL have the advantage of not requiring any time-modulated parameters (which may lead to undesired heating), while providing a large flux density, typically one or two flux quanta per unit cell [3]. However, OFL schemes proposed so far, when they are aimed at atomic species such as alkali-metal atoms, may induce significant photon scattering, which limits the available time during which the atomic gas remains cold enough to observe topology-induced effects [4, 5].
@@
In this paper, we propose a modified version of the OFL concept that operates with a so-called ‘dark state’, thus reducing considerably the heating problem due to photon scattering. Dark state lattices [6, 7] have recently been identified as a method for achieving sub-wavelength structures, either in 1D [8, 9, 10] or 2D [11, 12]. Here we propose a scheme providing an energy spectrum remarkably close to the Landau level structure for a charged particle in a uniform magnetic field. For realistic parameters, we obtain a series of many equidistant, non-overlapping, topological bands, each with a Chern index equal to 1. Being restricted to the dark internal state, the atoms remain weakly coupled to light, giving rise to delocalized orbitals. This behavior contrasts with other realizations of topological bands in optical lattices, which can be described by discrete lattice models such as the Haldane and Harper-Hofstadter models [13, 14, 15]. As an illustration of the topological character of the bands, we check that an ideal Fermi gas populating the ground band should exhibit an incompressible bulk surrounded by chiral edge states, and weakly interacting bosons should organize themselves into an ordered vortex lattice.
@@
We consider atoms for which a $\Lambda$-type optical transition can be isolated, with two states $|g_{\pm}\rangle$ from the ground-level manifold and a state $|e\rangle$ from an excited manifold (see Fig. 1a). The OFL is formed by two sets of monochromatic laser beams, with set 1 (resp. 2) at frequency $\omega_{1}$ (resp. $\omega_{2}$) close to the frequency $\omega_{+}$ of the $|g_{+}\rangle\leftrightarrow|e\rangle$ transition (resp. $\omega_{-}$ for $|g_{-}\rangle\leftrightarrow|e\rangle$) (see Fig. 1b). We assume that these frequencies fulfill the Raman resonance condition $\omega_{1}-\omega_{2}=\omega_{+}-\omega_{-}$ so that one can identify a space-dependent dark state, linear combination of $|g_{\pm}\rangle$ [7].
@@
Figure 1: (a) The $\Lambda$ level scheme considered in this manuscript, involving light fields at two close frequencies (red and blue arrows). (b) Spatial configuration of the lasers producing the dark optical lattice, with two standing waves along the directions $x\pm y$ for each frequency. (c) Band spectrum of the flux lattice as a function of the light coupling amplitude $V_{0}$. In the regime $V_{0}\gg\hbar\omega_{\text{c}}$, we find at low energy a succession of narrow energy bands with dominant dark-state character and an almost uniform spacing $\simeq\hbar\omega_{\text{c}}$. The red lines give a rough estimate of the first two bands with dominant bright-state character. They correspond to the energy levels $(n+1)\hbar\Omega_{\text{bright}}$ (with $n$ integer) in the lattice potential experienced by the bright state in the large depth limit of independent harmonic wells. (d) Mean value $\langle\hat{V}\rangle$ of the light shift as a function of the mean total energy $\langle E\rangle$, averaged over the quasi-momentum $\bm{q}$}, for the successive bands. The small value of the light shift for the low-energy bands shows that they predominantly populate the dark state.
The atom-laser coupling reads in the $\{|g_{+}\rangle,|g_{-}\rangle\}$ basis:
$\hat{V}(\bm{r})=V_{0}\begin{pmatrix}|\alpha_{+}|^{2}&\alpha_{+}^{*}\alpha_{-}\\ \alpha_{+}\alpha_{-}^{*}&|\alpha_{-}|^{2}\end{pmatrix}$, (1)
where $V_{0}$ is a real, positive energy and $\alpha_{\pm}(\bm{r})$ dimensionless complex amplitudes. Here we adiabatically eliminated the excited state $|e\rangle$ assuming that the detuning from resonance $\Delta\equiv\omega_{1}-\omega_{+}=\omega_{2}-\omega_{-}$ is positive and large compared to the Rabi frequency characterizing the atom-laser coupling (see appendix 1 for details) [16].
The dark state $|{\cal D}(\bm{r})\rangle\propto\alpha_{-}(\bm{r})|g_{+}\rangle-\alpha_{+}(\bm{r})|g_{-}\rangle$ is the eigenstate of $\hat{V}$ with energy 0. The other eigenstate of $\hat{V}$ is the bright state $|{\cal B}(\bm{r})\rangle\propto\alpha_{+}^{*}(\bm{r})|g_{+}\rangle+\alpha_{-}^{*}(\bm{r})|g_{-}\rangle$ with energy $E_{\text{bright}}({\bm{r}})=V_{0}(|\alpha_{+}({\bm{r}})|^{2}+|\alpha_{-}({\bm{r}})|^{2})$. This energy is positive everywhere so that the dark state is always the local ground state of $\hat{V}(\bm{r})$.
In this work, we will choose the following expressions for the coefficients $\alpha_{\pm}(\bm{r})$  [^1] An earlier version of this manuscript (arXiv:2412.15038v2) used a slightly more involved configuration with $\alpha_{+}=1+\frac{1}{2}\left[\cos(kx)+\cos(ky)\right]$ and $\alpha_{-}=\sin\left[\frac{k}{2}(x-y)\right]-{\rm i}\sin\left[\frac{k}{2}(x+y)\right]$. This configuration was subsequently explored in detail in [12].
$\alpha_{+}(\bm{r})$ $=$ $\sin X+{\rm i}\sin Y,$ (2)
$\alpha_{-}(\bm{r})$ $=$ $\cos X+\cos Y,$ (3)
where we set $X=k(x+y)/2$ and Y=k(yx)/2Y=k(y-x)/2. Here we assumed that the two wavenumbers of the beams forming the lattice, $k_{i}=\omega_{i}/c$ ($i=1,2)$, are close to each other and we set $k\sqrt{2}\equiv k_{1}\approx k_{2}$. The light-shift coupling $\hat{V}$ is then invariant upon discrete translations by $d\,{\bm{e}}_{x}$ and $d\,{\bm{e}}_{y}$ where $d=2\pi/k$, hence a square unit cell of area $d^{2}$.
The main results of this paper are based on the numerical calculation of the energy bands and of the associated Bloch states for a two-level atom moving in the potential $V(\bm{r})$. This will be done without any assumption regarding the adiabatic following of the dark state. However, it is interesting to start with the case when this adiabatic following is assumed. The adiabatic following of the dark state $|{\mathcal{D}}(\bm{r})\rangle$ gives rise to a synthetic magnetic field (the Berry curvature) [18], whose flux across the unit cell is quantized by an integer $N_{\phi}$ times the magnetic flux quantum $h$. The integer $N_{\phi}$ corresponds to the number of times the dark state wraps the Bloch sphere associated with the pseudo-spin $|g_{\pm}\rangle$ when moving across the unit cell [3]. Here, one can check that $N_{\phi}=1$, a non-zero value that characterizes optical flux lattices. For example, the north pole $|g_{+}\rangle$ is reached when ${\bm{r}}\to\bm{0}$, and the south pole $|g_{-}\rangle$ on the edges of the primitive unit cell ${\bm{r}}=\pm(d/2){\bm{e}}_{x}$ or $\pm(d/2){\bm{e}}_{y}$.
This non-zero flux is one of the requirements for an OFL. Another requirement formulated in [3] is that the eigenstates of the coupling $\hat{V}(\bm{r})$ should be non-degenerate at any point $\bm{r}$. This second requirement is clearly not fulfilled for the choice (2-3) since $\hat{V}$ vanishes at the unit cell corner ${\bm{r}}\to d({\bm{e}}_{x}+{\bm{e}}_{y})/2$. In fact, this second requirement cannot be satisfied for a dark-state based OFL [19]. As shown in appendix 2, the bright state admixture remains small at any point of the unit cell, so that the local vanishing of $\hat{V}$ does not prevent the occurence of topological bands.
Note that the configuration studied here yields the maximum possible flux density from light – about one flux quantum per $\lambda^{2}$, where $\lambda$ is the wavelength of the beams forming the lattice. This contrasts with early proposals for artificial gauge fields using dark states [20, 21] (see also [22] for a non-Abelian version). Those relied on configurations where $E_{\text{bright}}$ remained non-zero everywhere and assumed adiabatic following of $|{\cal D}(\bm{r})\rangle$, producing much lower flux densities – about one flux quantum per $w^{2}$ or $\lambda w$, with $w$ the laser beam size.
@@
We show in Fig. 1c the band spectrum for a two-level atom for the Hamiltonian $\hat{\bm{p}}^{2}/(2m)+\hat{V}(\bm{r})$ as a function of the light coupling amplitude $V_{0}$. Here, energies are expressed in units of the cyclotron frequency $\omega_{\text{c}}=2\pi\hbar/(md^{2})$ expected for a Landau level with the same magnetic flux density as our OFL. For large couplings $V_{0}\gg\hbar\omega_{\text{c}}$, we find a series of narrow energy bands at low energy, with an almost uniform spacing matching the cyclotron gap $\hbar\omega_{\text{c}}$. These bands essentially populate the dark state, with a very small residual light shift, shown in Fig. 1d for $V_{0}=100\,\hbar\omega_{\text{c}}$. For this light coupling, the ground band exhibits a width $\simeq 0.10\,\hbar\omega_{\text{c}}$ (about 8 times smaller than the gap to the first excited band), and the mean light shift is as low as $\langle\hat{V}\rangle\simeq 3\times 10^{-4}V_{0}$. The mixing between dark and bright states occurs at a higher energy, which can be estimated by computing the spectrum of an atom subjected to the lattice potential given by the bright-state energy $E_{\text{bright}}({\bm{r}})$. In the large $V_{0}$ limit, one can neglect tunneling between the different wells of the bright-state lattice, and one expects a harmonic spectrum for that state of frequency $\Omega_{\text{bright}}=\sqrt{2\pi V_{0}\omega_{\text{c}}/\hbar}$ (dashed red lines in Fig. 1c).
To characterize the topological nature of the bands, we compute their Chern number, and find the value $C=1$ for the first 19 bands at $V_{0}=100\,\hbar\omega_{\text{c}}$, confirming the similarity with Landau levels. Furthermore, the Berry curvature of the ground band is relatively uniform, with an rms deviation equal to $\simeq 12\%$ of its mean value. We study in [19] the quantum geometrical tensor, which can be used to characterize more precisely the analogy between the low-lying bands of our OFL and Landau levels.
A general comparison between OFL bands and ideal Chern bands is provided in [23], where it is shown that for the lattice described by Eqs.(2,3), the width of the ground band tends to 0 as $V_{0}/\hbar\omega_{c}\to\infty$. Furthermore, drawing on concepts from moiré materials, the authors of [23] demonstrate that, for a fixed $V_{0}$, this width can be reduced to extremely low values by introducing additional light beams, thereby reinforcing the proximity to the Landau level structure.
@@
We can extend the comparison between the OFL and Landau levels by inspecting individual Bloch states in real space. We first recall the Landau level formalism in a way that is well suited for comparison with flux lattices, i.e. a spatially periodic problem. We focus here on the ground band, the lowest Landau level (LLL). We assume a magnetic field ${\bm{B}}=B\,{\bm{e}}_{z}$ and adopt the Landau gauge ${\bm{A}}=-By\,{\bm{e}}_{x}$, for which the system is invariant upon $x$-translations $t_{x}(d_{x})=\exp(-d_{x}\,\partial_{x})$ of arbitrary distance $d_{x}$. The resulting conservation of the canonical momentum $p_{x}$ yields the LLL basis states $\lvert\phi_{p_{x}}\rangle$, of wavefunction
@@
where $\ell$ is the magnetic length. The basis $\lvert\phi_{p_{x}}\rangle$ is not convenient for comparison with OFL Bloch bands, which are indexed by a 2D quasi-momentum $\bm{q}$}. In order to symmetrize the $x$ and $y$ directions, we use another symmetry of the system, namely the $y$-magnetic translation symmetry $t_{y}(d_{y})=\exp[-d_{y}\,(\partial_{y}-{\rm i}x/\ell^{2})]$ [24]. Translation symmetries obey $t_{x}(d_{x})t_{y}(d_{y})={\rm e}^{-{\rm i}d_{x}d_{y}/\ell^{2}}t_{y}(d_{x})t_{x}(d_{y})$, hence a common basis of $x$- and $y$-translations can be found by choosing $d_{x}=d_{y}=d=\sqrt{2\pi}\ell$, such that the $d\times d$ unit cell is thread by a single unit of flux quantum $\Phi_{0}$. Magnetic Bloch states $\lvert\psi_{q_{x},q_{y}}\rangle$ invariant under both $t_{x}(d)$ and $t_{y}(d)$ can be formed by discrete sum of momentum states $\lvert\phi_{p_{x}}\rangle$, as
$\lvert\psi_{q_{x},q_{y}}\rangle=\sqrt{d}\sum_{n=-\infty}^{\infty}{\rm e}^{{\rm i}q_{y}dn}\lvert\phi_{q_{x}+nQ}\rangle$ (5)
with $Q=2\pi/d$. This expression corresponds to a discrete Fourier transform of the momentum states $\lvert\phi_{q_{x}+nQ}\rangle$ whose density probabilities are shifted one to the other along $y$ by integer multiples of $d$. The states $\lvert\psi_{q_{x},q_{y}}\rangle$, defined in the Brillouin zone $|q_{x}|,|q_{y}|<\pi/d$, form a basis of the LLL. We note that the wavefunctions match the LLL states defined on a torus of area $d\times d$ thread by magnetic fluxes $2\pi(\Phi_{x},\Phi_{y})/\Phi_{0}=(q_{x},q_{y})d$, given by [25]
$\psi_{q_{x},q_{y}}(x,y)=\frac{{\rm e}^{-y^{2}/(2\ell^{2})-{\rm i}q_{x}q_{y}d^{2}/(2\pi)}}{\pi^{1/4}\sqrt{\ell}}\Theta\begin{bmatrix}q_{x}d\\ q_{y}d\end{bmatrix}\left(\frac{x-{\rm i}y}{d}\middle|{\rm i}\right),$
@@
Figure 2: Mean density (a,c) and velocity (b,d) distributions on a square unit cell $d\times d$ thread by one unit of magnetic flux quantum. The left (resp. right) panels correspond to the lowest band of the OFL for $V_{0}=100\,\hbar\omega_{\text{c}}$ (resp. the LLL). In (b,d), the color encodes the norm of the velocity, and the red arrows the velocity streamlines. Similarly to the vortex singularity encountered in the LLL, the unit cell of the OFL exhibits a highly contrasted density dip around which the velocity circulate.
@@
is shown in Fig. 2d. We note that the circulation of velocity around each unit cell cancels due to the compensation between the singular contribution of the vortex and the rotational flow induced by the magnetic field. Each magnetic Bloch state thus exhibits non-zero local currents within the magnetic unit cell, but no global flow on larger length scales. The occurrence of vortices in OFL orbitals provides another illustration of the ‘ideal’ Chern band character: like in the LLL, OFL orbitals can be fully characterized by the location of vortices [27].
Figure 3: In situ density (a) and current density (b) profile of a Fermi gas at zero temperature in a quasi-2D geometry subjected to the optical flux lattice with $V_{0}=100\,\hbar\omega_{\text{c}}$ and to an isotropic harmonic trap (trapping frequency $\omega_{\text{trap}}=0.04\,\omega_{\text{c}}$). We used a chemical potential $\mu_{0}=1.79\,\hbar\omega_{\text{c}}$ located between the ground and first excited bands. A few stream lines are shown as red lines in (b). (c) Density profile along the line $y=0$, showing incompressibility in the bulk.